Determine the set of points at which the function is continuous
1) G(x,y)=tan^-1((x+y)^-2)
2) f(x,y,z)=√y÷(x^2-y^2+z^2)
3) f(x,y,z)=√z+y+z

4) f(x,y)={x^2y^3÷(2x^2+y^2 if (x,y)≠(0,0)
1 if (x,y)=(0,0)

5) f(x,y)={ xy÷(x^2+xy+y^2 if (x,y)≠(0,0)
0 if (x,y)=(0,0)
1
Expert's answer
2020-11-24T17:29:56-0500
  1. Now "tan^{-1}" is continuous function. Hence given function is continuous except when "x+y=0" in which case the function is not defined. Hence the function is continuous in its domain of definition i.e. "x+y\neq 0."
  2. In this function also, being a rational function is continuous every where in its domain of definition, namely, "y\geq 0, x^2-y^2+z^2\neq 0."
  3. This function is continuous at every point in its domain of definition, i.e. "z\geq 0."
  4. Let "x\neq 0." Let "y=mx." Then "f(x,mx)=\frac{m^3x^3}{2+m^2}" This function is always continuous. Now if "x=0," then "f(x,y)=0" if "y\neq 0." Now for any sequence "(x_n,y_n)\rightarrow (0,y), f(x_n,y_n)\rightarrow 0." Hence continuous. Now only point left is "(0,0)." Now if we take, "x=y," we get, "f(x,y)=\frac{x^3}{3}." So "lim _{x\rightarrow 0} f(x,x)=0." But "f(0,0)=1." Hence function is continuous everywhere except at (0,0).
  5. As in 4, "f(x,mx)=\frac{m}{1+m+m^2}". Hence continuous for "x\neq 0." Now if "x=0," then the function takes value 0. So for "(0,y), y\neq 0," "f(x,y)=0." Also for any sequence "(x_n,y_n)\rightarrow (0,y)," "f(x_n,y_n)""\rightarrow 0" since numerator tends to 0, but denominator is non-zero. Hence continuous at all points except may be at (0,0). Now taking "y=x=1\/n," we see that "f(x,y)=\frac{1}{3}" . Hence this doesn't converge to "f(0,0)=0" and hence not continuous at (0,0).